find an equation of th line that contains the given pair of points (13,21) (1,9)
Answer by Alan3354(69443)

find an equation of th line that contains the given pair of points (13,21) (1,9)
-------------
There's only one equation.
1st, find the slope, m
m = diffy/diffx = 12/12
m = 1
-----------
Use y = mx+b with either point to find b
9 = 1*1 + b
b = 8
--> y = x + 8